OpenStudy (asylum15):

solve for theta :

OpenStudy (asylum15):

\[\tan ^{2}\theta-13 = 5\sec \theta\] 0 degrees less than or equal to theta less than or equal to 360 degrees

OpenStudy (mertsj):

\[\frac{\sin ^2x}{\cos ^2x}-\frac{5}{\cos x}-13=0\] \[\sin ^2x-5\cos x-13\cos ^2x=0\] \[(1-\cos ^2x)-5\cos x-13\cos ^2x=0\] \[-14\cos ^2x-5\cos x+1=0\] \[14\cos ^2x+5\cos x-1=0\] \[(7\cos x-1)(2\cos x+1)=0\]

OpenStudy (mertsj):

Set each factor equal to 0 and solve.

OpenStudy (asylum15):

Thank you :)

OpenStudy (mertsj):

yw